OpenStudy (anonymous):

Solve 4/9 = -2v + 1/3 for V .

OpenStudy (anonymous):

solve for v what do you think your first step will be?

OpenStudy (anonymous):

i have no idea

OpenStudy (anonymous):

We can rearrange this and simplify so... 4/9 = -2v + 1/3 now 2v+1/9=0 next subtract 1/9 so 2v=-1/9 so v=-1/18 Do you follow that?

OpenStudy (anonymous):

yes, thanks!

OpenStudy (anonymous):

meh, as the answer is already given, ill just write the overall process/ method the goal when solving for variables is to isolate the variable so that its on one side this is achieved by adding, subtracting, multiplying or dividing so something like x+1=5 you want to get rid of the 1 on the left side, this can be done by subtracting 1 on the left side but in order to keep it equal, you need to do it to both sides you will then get x+1-1=5-1 1-1=0 so you get x=4 another example, but this time using division x/4=2 you want to get rid of the 4 in the denominator this can be done by multiplyying by 4 to both sides 4 x/4=2*4 4/4=1 x=8 remember, when you want to do something to one side, you must do it to the other if you dont, you'll get the wrong answer because the equation will not be balanced